# Uncertainty-Guided Optimization on Large Language Model Search Trees

### v2

@@ C.6 Full results for decoding experiments with LLama-2-7b and GPT-2
C.7 Details for experiment with repetition penalty
@@
Since we ULTS optimizes the total likelihood of a sequence and not the average likelihood of a sequence, we additionally scale the term with the tree depth, leading to $b(x)=d\cdot\log\textbf{diversity}((a_{1},...,a_{i}))$ in log-space. For ULTS we use a Dirichlet prior with concentration parameter $\alpha=5\times 10^{-6}$, $\epsilon=0.1$, $k_{max}=5$. and penalty parameter $\lambda\in\{0.0,0.2,0.4,0.6,0.8,1.0\}$. The parameters for contrastive search are top-$k$ $=5$ and penalty parameter $\alpha=0.6$. For beam search, we use $k=5$. The main paper reports results for ULTS with $\lambda=0.0$, i.w. without penalty, and $\lambda=0.2$. The full results are given in Table 3. One can see that with increasing penalty the diversity of the decoded sequences increases, i.e. extending the acquisition function with a repetition term is effective.
Method Perplexity $\downarrow$ Diversity $\uparrow$
Beam search 1.16$\pm$0.01 0.33$\pm$0.01
Nucleus sampling 3.69$\pm$0.07 0.82$\pm$0.01
Contrastive search 1.27$\pm$0.01 0.4$\pm$0.01
ULTS 1.16$\pm$0.01 0.33$\pm$0.01
ULTS with $\lambda=0.2$ 1.3$\pm$0.01 0.55$\pm$0.01
ULTS with $\lambda=0.4$ 1.45$\pm$0.01 0.74$\pm$0.01
ULTS with $\lambda=0.6$ 1.5$\pm$0.01 0.79$\pm$0.01
ULTS with $\lambda=0.8$ 1.53$\pm$0.01 0.82$\pm$0.01
ULTS with $\lambda=1.0$ 1.55$\pm$0.01 0.83$\pm$0.01

### v3

@@ C.8 Full results for decoding experiments with LLama-2-7b and GPT-2
C.9 Experiment with alternative utility function
In the following experiment, we evaluate ULTS with a different utility function – one that additionally contains a repetition-penalty term:
$\tilde{u}(x_{c},\boldsymbol{v}_{i})=\mathbb{I}[(\log v_{\hat{x}_{c}}+\lambda\,b(\hat{x}_{c}))=\max_{x_{j}\in\mathrm{children}(x_{i})}(\log v_{\hat{x}_{j}}+\lambda\,b(\hat{x}_{j}))],$ (7)
where $b(\hat{x}_{c})$ is a log-diversity term [54]. In this task, we complete 100 sentences from the Wikipedia dataset with GPT-2. We predict 100 tokens. We define the diversity term $b(x)$ of a node $x$ based on the proportion of duplicated $n$-grams in the token sequence $(a_{1},...,a_{i})$ corresponding to the node. For a token sequence $(a_{1},...,a_{i})$, let $\textbf{rep-n}((a_{1},...,a_{i}))=\bigl{(}1-\frac{|\text{unique $n$-grams}((a_{1},...,a_{i}))|}{|\text{total $n$-grams}((a_{1},...,a_{i}))|}\bigr{)}$. Following [54], we measure diversity by taking repetitions at different $n$-gram levels into account:
Since we ULTS optimizes the total likelihood of a sequence and not the average likelihood of a sequence, we additionally scale the term with the tree depth, leading to $b(x)=d\cdot\log\textbf{diversity}((a_{1},...,a_{i}))$ in log-space. For ULTS we use a Dirichlet prior with concentration parameter $\alpha=5\times 10^{-6}$, $\epsilon=0.1$, $k_{max}=5$. and penalty parameter $\lambda\in\{0.0,0.2,0.4,0.6,0.8,1.0\}$. The parameters for contrastive search are top-$k$ =5=5 and penalty parameter $\alpha=0.6$. For beam search, we use $k=5$. The main paper reports results for ULTS with $\lambda=0.0$, i.w. without penalty, and $\lambda=0.2$. The full results are given in Table 3. One can see that with increasing penalty the diversity of the decoded sequences increases, i.e. extending the acquisition function with a repetition term is effective.
Method Perplexity $\downarrow$ Diversity $\uparrow$
@@ C.9 Experiment with alternative utility function
[…]
@@
C.10 Additional metrics
Figure 14: ROUGE scores for the summarization task.
Figure 14 shows an evaluation in terms of ROUGE metrics for the summarization task on the TL;DR dataset described in Section 5. ROUGE 1 quantifies the overlap of unigrams between the produced sentence and ground truth reference, ROUGE 2 refers to bigrams, and ROUGE L refers to the longest commen subsequence. Though there is no perfect correlation between the ROUGE metric and the log-likelihood, ULTS still tends to achieve a good trade-off between performance and node expansions overall. Figure 15 shows additional results in terms of the BLEURT metric. Here, ULTS performs well for small $k_{max}$ but shows a decrease in performance for larger $k_{max}$.
Figure 15: BLEURT scores for the machine translation task.
C.11 ULTS with larger budget
Figure 16 shows additional results for ULTS with the empirical prior and GPT-2 on the Wikipedia dataset, as well as ULTS with the Dirichlet prior and t5 on the machine translation task (Figure 17).
Figure 16: Comparison of ULTS with the baselines on the Wikipedia dataset as described in Section 5 with additional data points for $k_{\max}\in\{50,100,200\}$.
Figure 17: Comparison of ULTS in the machine translation task as described in Section 5 with additional data points for $k_{\max}\in\{50,100\}$.
Appendix D Pseudocode
0: Number of tokens to generate $D$, number of samples $N$, approximate priors $\{\mathcal{B}_{l}\}_{l=1}^{D}$, confidence parameter $\varepsilon$, maximum number of expandable nodes per level $k_{\text{max}}$
0: A leaf $s^{*}$ and its likelihood $c^{*}$
{Initialization}
[…]
2:for $n\leftarrow 1$ to $N$ do
3:  Sample $\Delta_{s_{0},n}\sim\mathcal{B}_{1}(\Delta)$
4:  $v_{s_{0},n}\leftarrow\log\Delta_{s_{0},n}$
5:end for
[…]
7:while $\hat{\mathbb{P}}(c^{*}<v_{s_{0}})>\varepsilon$ do
8:  //Selection always starts from root
9:  $s_{i}\leftarrow\texttt{select}(s_{0},k_{\text{max}})$
//Expand
10:  $\mathcal{L}\leftarrow(\mathcal{L}\setminus\{s_{i}\})\cup\texttt{children}(s_{i})$
11:  for $s_{c}\in\texttt{children}(s_{i})$ do
12:   //Generate posterior samples
13:   for $n\leftarrow 1$ to $N$ do
14:    Sample $\Delta_{s_{c},n}\sim\mathcal{B}_{\texttt{level}(s_{c})}(\Delta)$
15:    $v_{s_{c},n}\leftarrow\log c_{s_{0}\to s_{c}}+\log\Delta_{s_{c},n}$
16:   end for
17:   //Update best complete path so far
18:   if $\texttt{level}(s_{c})=D$ $\log c_{s_{0}\to s_{c}}>c^{*}$ then
19:    $c^{*}\leftarrow\log c_{s_{0}\to s_{c}}$
20:    $s^{*}\leftarrow s_{c}$
21:   end if
22:  end for
23:  $\texttt{backup}(\{v_{s_{c},n}\}_{n=1}^{N},s_{0}\to s_{c})$
24:  //Termination probability
25:  $\hat{\mathbb{P}}(c^{*}<v_{s_{0}})\leftarrow\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}[c^{*}\leq v_{s_{0},n}]$
26:end while
Algorithm 2 ULTS